OpenStudy (anonymous):

HELP!! A line has a slope of 3 and contains the point (−1, −8). Which equations represent the line? Choose all answers that are correct. A. −3x + y = −5 B. 3x + y = −5 C. (y – 8) = 3(x − 1) D. (y + 8) = 3(x + 1)

jimthompson5910 (jim_thompson5910):

what is the equation in point slope form?

OpenStudy (anonymous):

Y=mx+b

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -1}}\quad ,&{\color{blue}{ -8}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= 3 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

is it B and D

OpenStudy (jdoe0001):

well.. what did you get?

jimthompson5910 (jim_thompson5910):

Y=mx+b is slope intercept form

OpenStudy (anonymous):

YEs it think it is B and D

OpenStudy (anonymous):

Can you help me with another one

jimthompson5910 (jim_thompson5910):

if you think it is B, plug (x,y) = (-1,-8) into equation B and test it out

jimthompson5910 (jim_thompson5910):

what do you get when you do that?

OpenStudy (anonymous):

A line has a slope of -1/2 and a y-intercept of 3. Which equation repsent the line? choose all answers. A. 2y=6+x B. x+2y=6 C. y= -1/2x+3 D. (y-2)=-1/2(x-2)